OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of two x plus three, over five.

OpenStudy (johnweldon1993):

\[f(x) = \frac{2x + 3}{5}\] Inverse = switch the 'x' and the 'y' and solve for 'y' again...so \[x = \frac{2y + 3}{5}\] Multiply both sides by 5 \[5x = 2y + 3 \] Subtract 3 from both sides \[5x - 3 = 2y\] Divide everything by 2 to isolate and solve for y \[y = \frac{5x - 3}{2}\] Make sense?

OpenStudy (anonymous):

yess , thankyou